OpenStudy (miasmith):

1.Given the system of constraints name all vertices of the feasible region. then find the maximum value of the given objective function. Show work constraints: x≥0 y≥0 y≤1/3x+3 5≥y+x objective function: C=6x-4y

OpenStudy (mathmale):

Graph the following LINES: x=0, y=0, y=(1/3)x+3, and y=-x+5. That would get you started. Identify the coordinates of every intersection of these lines. For example, the point of intersection of the line y=-x+5 with the vertical axis (x=0) is (0,5), which we call the "vertical intercept."